# Multi-Hypothesis Prediction for Portfolio Optimization: A Structured Ensemble Learning Approach to Risk Diversification

### v4

@@ 3.3 Multi-Hypothesis Portfolio Optimization with Structured Radial Basis Function Network (s-RBFN)
The optimal s-RBFN parameters are obtained by least-squares with the regularization parameter $\lambda_{s}$ and target portfolio $\boldsymbol{r}$:
$\boldsymbol{w}=\left(\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)^{\rm T}\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)+\lambda_{s}\ast\boldsymbol{I}_{(mxm)}\right)^{-1}\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)^{\rm T}\boldsymbol{r}$ (9)
The portfolio weights are the scaled s-RBFN parameters, which can be scaled in various ways, ensuring they sum to one and remain each within the range of 0 to 1. This particular scaling reflects the constraints of long-only portfolios, satisfying $\mathbf{1}^{\top}\boldsymbol{w}=1$ and $0\leq\boldsymbol{w}\leq 1$. After the optimization in (9), the target portfolio time-series returns would be given by:
$\boldsymbol{r}=\boldsymbol{w}_{\text{opt}}^{\top}\boldsymbol{x}$ (10)
with $\boldsymbol{w}_{\text{opt}}$ being the optimal weights and $\boldsymbol{x}$ the time series of the portfolio constituents. For $\boldsymbol{w}$ to be optimal in equation (9) $\boldsymbol{r}$ must be given by (10). To ensure in the PSEM that optimization is guided by the structure of the hypothesis prediction space of the constituents and the degree of diversification within the structured dataset of predictions already incorporated into $\boldsymbol{P}$, a canonical basis of the hypotheses is selected. where $\boldsymbol{P}=\left(\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)^{\top}\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)+\lambda_{s}\ast\boldsymbol{I}_{(m\times m)}\right)^{-1}\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)^{\top}$ consequently reducing (9):
$\boldsymbol{w}=\boldsymbol{P}\boldsymbol{r}$ (11)
The weights in equation (11) are selected such that the target portfolio $\boldsymbol{r}$ is equal-weighted, $\boldsymbol{w}=\left\{\frac{1}{\|\boldsymbol{x}\|}\right\}$. This ensures that the optimization and final solution for the portfolio weights are based on the information in $\boldsymbol{P}$ and not on other pre-specified targets:
$\boldsymbol{w}_{\text{opt}}=\boldsymbol{P}\frac{\boldsymbol{x}}{\|\boldsymbol{x}\|}$ (12)
with $\frac{\boldsymbol{x}}{\|\boldsymbol{x}\|}$ forming the canonical basis of hypotheses. Additionally, the centroid combiner rule in the bias-variance-diversity decomposition for the case of the MSE loss in (2), when the ensemble is a weighted combination as in the s-RBFN case, is given by $\boldsymbol{\bar{r}}=\frac{1}{M}\sum_{j=1}^{M}w_{j}\phi(\boldsymbol{\hat{x}}_{j}(\varepsilon))$, which is equivalent to the equal-weighted portfolio $\frac{1}{M}\sum_{j=1}^{M}\boldsymbol{x}_{j}$, if the assets are taken as the hypotheses instead of the weighted basis functions.
In Figure 2, a comparative analysis between two portfolio optimization frameworks is presented: the classical mean-variance approach from MPT (Markowitz, 1952), shown on the left, and the s-RBFN optimized via least squares (Dominguez et al., 2025), shown on the right. Despite their structural differences, both methodologies exhibit notable conceptual and computational parallels.
@@
Both approaches involve matrix multiplication as a core computational operation in portfolio optimization, rendering them computationally equivalent in this respect. In the PSEM framework using the s-RBFN model, the input data $\boldsymbol{x}$ is replaced with the diversified structured data set of predictions $\boldsymbol{\hat{x}}(\varepsilon)$, which is augmented through the use of basis functions. These basis functions—such as radial, Gaussian, or other types—introduce additional flexibility to the s-RBFN model, enabling it to capture more complex non-linear relationships within the data.
Figure 2: Systems representation comparison of the Mean-variance framework to portfolio optimization from Modern Portfolio Theory (Markowitz, 1952) (Left) versus the PSEM framework with the s-RBFN optimized by least squares (Dominguez et al., 2025) (Right). The portfolio constituents’ returns, $\boldsymbol{x}$, and the target portfolio returns, $\boldsymbol{r}$, in the left case are assumed to be normally distributed. The remaining parameters and variables include the portfolio weights $\boldsymbol{w}$, the diversity parameter $10<\varepsilon<$ 1 used for the structured data set $\boldsymbol{\hat{x}}(\varepsilon)$ formation, the individual predictors $\{f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{j})\}_{j=1}^{M}$, and the s-RBFN model basis functions in matrix form $\boldsymbol{\Phi}$.
The mean-variance approach is a fully supervised method, where the portfolio optimization process is guided by a predefined target return series. In contrast, the PSEM framework operates within a semi-supervised learning paradigm, as it does not rely on a predefined target. Instead, the optimal portfolio weights are determined based on the hypotheses generated by the predictors and their mixtures during training. Consequently, to provide a canonical reference for the hypothesis space, it is necessary to employ an equal-weighted portfolio as a target portfolio ($\boldsymbol{r}=\frac{1}{||\boldsymbol{x}||}\boldsymbol{x}$ in Figure 2). This ensures that the decision-making is governed solely by the structured data set $\boldsymbol{\hat{x}}(\varepsilon)$ and the s-RBFN model, thereby avoiding any dependence on a specified target.
The next section describes the methodology used to conduct portfolio allocation experiments, encompassing both the asset selection and weight optimization stages, from the perspective of multi-hypothesis prediction with structured models. This approach seeks to uncover connections between diversity in asset selection and out-of-sample portfolio diversification, identify new limits on portfolio diversification introduced by this additional source of variability, and evaluate the ability to control these effects parametrically prior to weight optimization.
@@ 3.4 Parametric Control of Out-of-Sample Diversification at the Asset or Hypothesis Selection Stage
In conclusion, modeling the portfolio allocation problem as a structured prediction task over multiple hypotheses allows the separation of the process into two key components: Hypotheses selection and weight optimization. This approach provides parametric methodologies to induce diversity in hypothesis selection through the parameter $\gamma$, and to regulate diversity in the learning process of individual predictors based on the selected hypotheses through the parameter $\varepsilon$. These two sources of diversity are directly connected to out-of-sample portfolio risk diversification and can be controlled and adjusted prior to the optimization and decision-making process.
@@ 4 Numerical Results
4.1 Data and Setup
The time series of daily prices for the S&P 500 index members (S&P 500, 2005) over the past six years (excluding weekends) was obtained from Bloomberg (S&P 500, 2024). Returns were calculated as the percentage change in daily prices.
@@ 4.1 Data and Setup
The structured ensemble models used are the s-RBFN optimized by least squares and two-layer networks as individual predictors with the same tunned hyperparameters for all the cases as described in (Dominguez et al., 2025). The equal-weighted model and the MSE-weighted model are also included, where the weights are either equally distributed across all constituents or inversely proportional to their generalization prediction error. For asset selection, the ranking is constructed by calculating the one-month recurrent prediction of the daily time series and then ranking based on the accumulated monthly return. The prediction is performed using the same model as the individual predictors with identical hyperparameters, but each of the 500 index members has its own network.
@@ 4.2.1 Diversity in the Ensemble Learning Stage
In Figure 6, the same set of experiments but for the sequential multi-step decision-making case is presented. In this case, the resemblance is different from that shown in MPT and Figure 4, which makes sense since those setups are one-step, whereas the results here are sequential multi-step. Nevertheless, the conclusions remain unchanged, and the same patterns can be observed regarding the relationship between diversity in ensemble learning and out-of-sample portfolio diversification. A consistent increase in Sharpe ratios can be seen as the parameter $\varepsilon$ and the number of predictors, hypotheses, or assets $M$ increase. On the other hand, the regularization parameter has a greater impact in this case than in the one-step scenario shown in Figure 5. It can be observed that the out-of-sample risk-adjusted portfolio performance improves even in the case where the diversity parameter $\varepsilon$ is not applied. This may indicate that performance in multi-step settings depends not only on diversification but also on the complexity of the data. The same experiments can be seen for the case of radial basis functions with the same conclusions in the Appendix in Figures 9 and 10.
@@
4.2.2 Including Diversity in the Asset or Hypothesis Selection Stage
@@
Figures 7 and 8 present the out-of-sample Sharpe ratios for different numbers of stocks $M$, various values of the asset selection diversity parameter $\gamma$, and the learning stage diversity parameter $\varepsilon$, as well as threshold values $T$ used in the asset selection process based on forecast rankings, for the multi-step sequential (10-month strategy) case. Figures 12 and 13 in the Appendix show the same results for the one-step decision case (1-month, 100 simulations).
@@ 5 Conclusion and Future Work
This work shows that valuable concepts of modern portfolio theory
(Markowitz, 1952) that are the basis of modern finance can be understood through structured models for multiple hypotheses prediction, providing new insights. By framing the problem this way, the diversity of individual predictors can be linked with portfolio diversification, offering a fresh perspective based on learning in multiple hypothesis prediction settings. The proposed PSEM strategy enables a structured framework for portfolio optimization, using as portfolio target the arithmetic combiner or equal-weighted portfolio. To the authors’ knowledge, the first plug-in method that parametrically controls out-of-sample diversification prior to the decision-making is introduced. This is useful for practitioners, as they can access many diversified versions of the structured data set with different levels of diversification before portfolio optimization.
This work confirms that the out-of-sample diversification and generalization performance of a portfolio depends not only on the optimization process but also on asset selection. To the authors’ knowledge, this is the first time a method that parametrically induces out-of-sample portfolio diversification from the asset selection stage is presented in the literature. Experiments validate the diversity-quality trade-off of return predictions in asset selection discussed in Section 3.4. The experiments show that, for both the one-step and multi-step cases, when the RMSE ranges in prediction are similar, asset selections with greater diversity in the predicted values prevail over less diverse selections, even if the latter have more positive average returns as a group.
Surprisingly, the experiments demonstrate that the out-of-sample diversification can be increased through diversity in asset selection, even with a constant number of constituents (in contrast to diversification limits (Dalio, 2018). This is probably due to the fact that other frameworks do not take into account the diversification component resulting from asset selection diversity. This is particularly useful for practitioners with constraints on the number of assets who aim to diversify their portfolios. The proposed approach offers parametric methods to introduce diversity in the selection of hypotheses via the parameter $\gamma$ and to manage diversity in the training process of individual predictors based on these hypotheses through the parameter $\varepsilon$. Both sources of diversity are closely linked to out-of-sample portfolio risk diversification and can be adjusted prior to the optimization and decision-making stages, which is particularly useful for practitioners.
For future work, it would be interesting to investigate the application of deeper architectures for individual predictors and multimodal datasets. This would enable the construction of portfolios where the hypotheses are not limited to stock time-series but can include any source of information relevant to any particular hypothesis. Additionally, it would be valuable to test the proposed methodology in scenarios where the hypothesis is a group of stocks or a portfolio, and to explore how to choose an appropriate target in such cases. Another potential direction for future research is to examine the relationship between complexity and diversity in the ensemble loss within the PSEM framework. Finally, an interesting financial application would be the development of a risk premia factor that takes long positions in high-diversity portfolios and short positions in low-diversity portfolios, aiming to profit from the diversity-quality trade-off in predicted returns presented in this work.

### v5

@@ 3.4 Example with Structured Radial Basis Function Network (s-RBFN)
The optimal s-RBFN parameters are obtained by least-squares with the regularization parameter $\lambda_{s}$ and target portfolio, the arithmetic combiner or equal-weighted portfolio, $\boldsymbol{\bar{r}}$:
$\boldsymbol{w}=\left(\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)^{\rm T}\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)+\lambda_{s}\ast\boldsymbol{I}_{(mxm)}\right)^{-1}\boldsymbol{\Phi}\left(\boldsymbol{\hat{x}}(\varepsilon)\right)^{\rm T}\boldsymbol{\bar{r}}$ (10)
The optimal portfolio weights are equivalent to the s-RBFN optimal parameters $\boldsymbol{w}$, after applying weight normalization and other relevant constraints.
In Figure 2, a comparative analysis between two portfolio optimization frameworks is presented: the classical mean-variance approach from Modern Portfolio Theory (MPT) (Markowitz, 1952), shown in the blue box on the left, and the s-RBFN model trained via least squares (Dominguez et al., 2025), as the PSEM model, shown in the blue box on the right. Despite their structural differences, both methodologies exhibit notable conceptual and computational parallels.
@@
Both approaches allow closed-form solutions via least squares, rendering them computationally equivalent in terms of efficiency. In the PSEM framework, the input data $\boldsymbol{x}$ is replaced with the diversified structured dataset of predictions $\boldsymbol{\hat{x}}(\varepsilon)$, which is then augmented using basis functions within the s-RBFN model. These basis functions—such as radial, Gaussian, or other types—introduce additional flexibility to the s-RBFN model, enabling it to capture more complex non-linear relationships within the data.
Figure 2: Systems representation comparison between the Mean-Variance framework for portfolio optimization from Modern Portfolio Theory (Markowitz, 1952) (left blue box) and the predict-then-optimize PSEM framework (right blue box). In the Mean-Variance case, the portfolio constituents’ returns, $\boldsymbol{x}$, and the target portfolio returns, $\boldsymbol{r}$, are assumed to be normally distributed. In contrast, in the PSEM framework, the target is the equal-weighted portfolio $\boldsymbol{\bar{r}}$. Additional parameters and variables include the portfolio weights $\boldsymbol{w}$, the diversity parameter $10<\varepsilon<$ 1 used for forming the structured dataset $\boldsymbol{\hat{x}}(\varepsilon)$, the individual predictors $\{f_{\boldsymbol{\theta}_{j}}(\boldsymbol{x}_{j})\}_{j=1}^{M}$, and the s-RBFN model basis functions represented in matrix form by $\boldsymbol{\Phi}$.
The focus of this work is on portfolio allocation, which encompasses both asset selection and portfolio optimization. The latter has been addressed in the previous sections. The former is introduced in the next section through a novel method that leverages the diversity of asset return predictions as a asset selection tool. This approach seeks to uncover connections between diversity in asset selection and out-of-sample portfolio diversification, identify new limits on portfolio diversification introduced by this additional source of variability, and evaluate the ability to control these effects parametrically prior to weight optimization.
@@ 3.5 Parametric Control of Out-of-Sample Diversification at the Asset or Hypothesis Selection Stage
In conclusion, modeling the portfolio allocation problem as a structured prediction task over multiple hypotheses allows the separation of the process into two key components: Hypotheses selection and weight optimization. This approach provides parametric methodologies to induce diversity in hypothesis selection through the parameter $\gamma$ and to regulate diversity in the learning process of individual predictors based on the selected hypotheses through the parameter $\varepsilon$. These two sources of diversity are directly connected to out-of-sample portfolio risk diversification and can be controlled and adjusted prior to the optimization and decision-making process.
@@ 4 Numerical Results
4.1 Data and Experimental Setup
The time series of daily prices for the S&P 500 index members (S&P 500, 2005) (excluding weekends) was obtained from Bloomberg (S&P 500, 2024). Returns were calculated as the percentage change in daily prices.
@@ 4.1 Data and Experimental Setup
The structured ensemble models used are the s-RBFN optimized by least squares and two-layer networks as individual predictors with the same set of hyperparameters configurations for all the cases as described in (Dominguez et al., 2025). The equal-weighted model and the MSE-weighted model are also included, where the weights are either equally distributed across all constituents or inversely proportional to their generalization prediction error. For asset selection, the ranking is constructed by calculating the one-month recurrent prediction of the daily time series and then ranking based on the accumulated monthly return. The prediction is performed using the same model as the individual predictors with identical hyperparameters, but each of the 500 index members has its own network.
4.2 Model Evaluation and Performance Metrics
One of the most commonly used portfolio performance metrics is the Sharpe Ratio (Sharpe, 1994), defined as the ratio of the average daily return over a period to its standard deviation. In this work, return predictions are evaluated using 1-month cumulative returns. Accordingly, a slight variation of the Sharpe Ratio is considered, where the cumulative return over one month is divided by the standard deviation of daily returns during that same period. This modified Sharpe Ratio is used to assess out-of-sample portfolio performance in the one-step decision scenario. For the multi-step decision setting, the modified Sharpe Ratio is computed over a 10-month sequential investment horizon.
For the individual predictors, a two-layer multi-layer perceptron (MLP) is used with hyperbolic tangent activation functions. The number of neurons in each layer is denoted by $\kappa$, the learning rate by $\eta$, and a multiplicative factor for the random initial weights during gradient descent by $\chi$, which affects model performance. Regularization parameters are represented by $\lambda_{p}$. These individual predictors are applied during the asset selection stage, where RMSE is used as the time-series forecasting performance metric. Assets are selected from a pool of diversified assets based on their predicted values, with diversity scaled by the parameter $\gamma$.
The s-RBFN is employed as the PSEM model in the experiments, where the number of predictors or hypotheses is denoted by $M$, and the regularization parameter is represented by $\lambda_{s}$. Once the optimal portfolio weights are obtained—by normalizing the ensemble parameters—they are used to evaluate out-of-sample portfolio performance using the modified Sharpe Ratio: 1-month for the one-step decision case and 10-month for the multi-step decision case. Portfolio performance serves as the criterion for selecting the optimal network configuration for both the individual predictors and the s-RBFN. Multiple hyperparameter combinations are tested in the experiments, and all hyperparameter values used are listed in Table 1(b).
Table 1: Sets of values for the individual predictors and s-RBFN hyperparameters, (Dominguez et al., 2025).
(a) $M$ number of hypotheses, $\kappa$ number of neurons per layer, $\eta$ learning rates for the predictors, $\chi$ is a multiplicative factor for random initial predictors’ parameters $\Theta$.
$M$ $\kappa$ $\eta$ $\chi$
$[2,5,10,20,35]$ $[20,200,2000]$ $[0.03,0.3]$ $[0.0001,0.01,0.1,1]$
(b) $\varepsilon$ is the diversity parameter in the optimization and $\gamma$ in the asset selection, $\lambda_{p}$ is the regularization parameter for the predictors and $\lambda_{s}$ for the s-RBFN.
$\varepsilon$ $\lambda_{p}$ $\lambda_{s}$ $\gamma$
$[0,0.1,0.35,0.5]$ $[0,0.0001,0.01,0.07]$ $[0,3,5]$ $[1,2,3,5]$
@@ 4.3.1 Diversity in the Ensemble Learning Stage
In Figure 6, the same set of experiments but for the sequential multi-step decision-making case is presented. In this case, the resemblance is different from that shown in MPT and Figure 4, which makes sense since those setups are one-step, whereas the results here are sequential multi-step. Nevertheless, the conclusions remain unchanged, and the same patterns can be observed regarding the relationship between diversity in ensemble learning and out-of-sample portfolio diversification. A consistent increase in Sharpe ratios can be seen as the parameter $\varepsilon$ and the number of predictors, hypotheses, or assets $M$ increase. On the other hand, the regularization parameter has a greater impact in this case than in the one-step scenario shown in Figure 5. It can be observed that the out-of-sample risk-adjusted portfolio performance improves even in the case where the diversity parameter $\varepsilon$ is not applied. This may indicate that performance in multi-step settings depends not only on diversification but also on the complexity of the data. The same experiments can be seen for the case of radial basis functions with the same conclusions in the Appendix in Figures 10 and 11.
@@
4.3.2 Diversity-quality Trade-off of Returns Predictions: Including Diversity in the Asset Selection Stage
@@
Figures 7 and 8 present the out-of-sample Sharpe ratios for different numbers of stocks $M$, various values of the asset selection diversity parameter $\gamma$, and the learning stage diversity parameter $\varepsilon$, as well as threshold values $T$ used in the asset selection process based on forecast rankings, for the multi-step sequential (10-month strategy) case. Figures 13 and 14 in the Appendix show the same results for the one-step decision case (1-month, 100 simulations).
@@
4.4 Comparative Performance Analysis
This section presents a comparative analysis between the proposed method, implemented with the s-RBFN as the PSEM model, and several well-established robust and data-driven ”plug-in” portfolio optimization methods, including Inverse Volatility (IV) (ANG et al., 2006), CVaR Risk Parity (CVaR RP) (Kapsos et al., 2018), Maximum Diversification (MD) (Choueifaty and Coignard, 2008), Hierarchical Risk Parity (HRP) (Lopez de Prado, 2016) and Hierarchical Equal Risk Contribution (HERC) (Raffinot, 2018). The comparison is conducted over a two-year investment horizon with monthly portfolio reallocation. The performance metric used is the monthly standard Sharpe Ratio, calculated using a time series of 1-month cumulative returns. This is computed as the average return over the 24-month period divided by its standard deviation and then annualized. It is important to note that the Sharpe ratios reported in this section are generally lower because they are calculated using average returns, whereas in previous sections a modified version of the Sharpe ratio was used, based on cumulative returns over 1-month and 10-month periods.
Table 4 reports the average Sharpe Ratios across various threshold levels $T$ and values of the asset selection diversity parameter $\gamma$. It can be shown how the presented method to select assets based predictions diversity (diversify-quality trade-off of predictions as in Section 4.3.2) makes all the methods improve generalization performance due to increased out-of-sample diversification measured by the Sharpe ratios. Also, the s-RBFN method is the best as can be seen in the average value across all the experiments in the same table at the bottom. For selections of assets with more than -0.5% cumulative monthly return as in Table 4(a), diversity is not of clear value but for selections above 0.0% and 0.3% it is a clear wining selection choice for all methods.
Table 4: Average Sharpe ratios across all configurations of the optimization diversity parameter $\varepsilon\in\{0,0.1,0.35\}$ and number of portfolio constituents $M\in\{5,10,20,35\}$, grouped by threshold $T$ and the asset selection diversity parameter (multiplicative factor) $\gamma\in\{1,2,3\}$.
(a) Threshold $T=-0.5$
$\gamma$ s-RBFN IV CVaR RP MD HRP HERC
1 0.79 0.59 0.74 0.41 0.54 0.55
2 0.64 0.40 0.50 0.29 0.36 0.28
3 0.55 0.41 0.46 0.36 0.37 0.42
(b) Threshold $T=0.0$
$\gamma$ s-RBFN IV CVaR RP MD HRP HERC
1 0.70 0.40 0.59 0.31 0.36 0.39
2 0.85 0.56 0.63 0.55 0.55 0.41
3 0.69 0.43 0.49 0.29 0.43 0.51
(c) Threshold $T=0.3$
$\gamma$ s-RBFN IV CVaR RP MD HRP HERC
1 0.55 0.45 0.64 0.35 0.45 0.42
2 0.58 0.40 0.57 0.32 0.37 0.52
3 0.83 0.49 0.65 0.33 0.48 0.47
Avg. 0.69 0.46 0.58 0.35 0.44 0.44
Figure 9 displays the distribution of Sharpe ratios, ordered from highest to lowest, across all experiments conducted in this section (108 in total). These experiments include various configurations of the diversity parameters $\varepsilon$ and $\gamma$, threshold $T$, and the number of portfolio constituents $M$. The horizontal axis represents individual experiments, sorted by decreasing Sharpe ratio. The plot clearly demonstrates that the s-RBFN consistently outperforms all other models—often by a significant margin—highlighting the robustness and accuracy of the proposed method.
Figure 9: Sharpe ratios distributions across all experiments (108 configurations for $\varepsilon$,$\gamma$,$M$,$T$) for each portfolio optimization model.
The results indicate that the s-RBFN method consistently outperforms the other portfolio optimization approaches across all threshold levels and values of the asset selection diversity parameter ($\gamma$), achieving the highest average Sharpe ratios in most cases. Furthermore, performance generally improves for all methods as $\gamma$ increases, suggesting that greater diversity in asset selection positively contributes to out-of-sample performance.
@@ 5 Conclusion and Future Work
This work explored the relationship between the diversity of individual predictors in structured ensembles and portfolio diversification, offering a novel perspective within a multiple-hypothesis predict-then-optimize framework. One key contribution was the use of the ensemble combiner rule as the portfolio target during optimization—for example, the equal-weighted portfolio with the MSE loss. To the best of the authors’ knowledge, this represents the first plug-in method that enables parametric control of out-of-sample diversification prior to decision-making. Experimental results confirmed that out-of-sample diversification and generalization performance depend not only on the optimization procedure but also on the asset selection stage. Notably, when RMSE ranges were similar, asset selections with greater diversity in return predictions consistently outperformed those with less diversity, highlighting the importance of the diversity-quality trade-off. This forms a third contribution, and both sources of diversity were shown to extend the boundaries of portfolio diversification beyond previously suggested limits (Dalio, 2018).
For future work, it would be valuable to explore deeper architectures for individual predictors, as well as the integration of multimodal datasets, enabling hypotheses to incorporate information beyond stock time series. Furthermore, extending the methodology to settings where each hypothesis represents a portfolio or a group of assets could open new avenues for application. Another promising direction involves investigating the interplay between ensemble complexity, prediction diversity, and portfolio risk. Additionally, exploring alternative loss functions and ensemble combiners tailored to various portfolio risk objectives may further enhance the flexibility and effectiveness of the proposed framework.